OpenStudy (kirbykirby):

How does x^2 + y^2 = 4x give the equation of a circle?

OpenStudy (anonymous):

subtract 4x from both sides, then solve for y. this should give you the formulas for the upper and lower semicircle

OpenStudy (anonymous):

not exactly you'll have to complete the square

OpenStudy (anonymous):

well, yeah you can do that do

OpenStudy (anonymous):

if you subtract -4x you'll get 0=y^2+x^2-4x... if you complete the square you'll get \[y^2+x^2-4x+4=4\] \[=y^2+(x-2)^2=4\] you can go farther if you'd like but you don't need to divide by 4 and you'll get the 1 by itself which is considered another equation of a circle

OpenStudy (anonymous):

\[=\frac{y^2}{4}+\frac{(x-2)^2}{4}\]

OpenStudy (anonymous):

=1

OpenStudy (kirbykirby):

Is completing the square the only otion it seems?

OpenStudy (anonymous):

i believe so or you can do it the way that lagrange said but i think what they call his is Cartesian and it gives two semi circles which complete the circle

OpenStudy (kirbykirby):

Oh ok thanks!!